Calculate the pH expected for the 0.0100 M HCl solution used in part A.

Intermediate value Final value 1a. Calculate the pH expected for the 0.0100 M HCl solution used in part A. 1b. Calculate thePART D: INVESTIGATING PHOSPHATE BUFFERS Units value Value Desired pH unitless A 7 to 9 Acid chosen H2PO4 Kg of chosen acid 6.I'm so lost please help
Answer #1

ANSWER:

You have a 0.0100 M HCl solution. The HCl is a strong acid with the following chemical equation for the dissociation of the acid in water:

HCl\, (aq)\rightarrow H^{+}\, (aq)+Cl^{-}\, (aq)

Strong acids have a complete dissociation, this means that HCl is transformed completely into Cl- and H+. Then, 1 mol of HCl produces 1 mol of H+. Thus, in a solution of HCl:

[HCl]=[H^{+}]=0.0100 \, M

Then, for the solution of 0.0100 M HCl, the concentration of H+ would be 0.0100 M, and the expected pH for this solution is

pH = -log[H+] = -log(0.0100)

\mathbf{pH=2.00}

The percent error between expected pH and measured pH is calculated with this equation:

\%\, error=\frac{\left | pH_{measured}-pH_{expected} \right |}{pH_{expected}}\times 100\, \%

\%\, error=\frac{\left | 2.00-2.00 \right |}{2.00}\times 100\, \%

\mathbf{\%\, error=0.00\, \%}